The compound known as butylated hydroxytoluene, abbreviated as BHT, contains carbon, hydrogen, and oxygen. A 3.876 g sample of BHT was combusted in an oxygen rich environment to produce 11.61 g of CO2(g) and 3.803 g of H2O(g). Insert subscripts below to appropriately display the empirical formula of BHT.
1
Expert's answer
2018-02-15T13:55:08-0500
n(CO2)=m/M=11.61/44.01=0.2632 mol
n(H2O)=m/M=3.0803/18.02=0.1709 mol
m(C)=m*M=0.2638*12.01=3.1682 g
m(H)=m*M=0.1709 *2*1.008=0,3446g
m(O)=3.879-3.1682-0.3446=0.3662 g
n(O)= 0.3662/16.00=0.0227 mol
n(C):n(H):n(O)= 0.2638:2*0.1709:0.0227=12:15:1
Empirical formula C12H15O